A human cannonball at the circus is launched from a cannon at 45 m/s and an angle of 55° from the ground. Calculate the x and y components of the initial velocity.
1
Expert's answer
2020-11-17T11:31:01-0500

Let's write the "x"- and "y"-components of the initial velocity:


"v_x=vcos\theta=45\ \dfrac{m}{s}\cdot cos55^{\circ}=25.81\ \dfrac{m}{s},""v_y=vsin\theta=45\ \dfrac{m}{s}\cdot sin55^{\circ}=36.86\ \dfrac{m}{s}."

Answer:

"v_x=25.81\ \dfrac{m}{s}."

"v_y=36.86\ \dfrac{m}{s}."